# A NIRCam-dark galaxy detected with the MIRI/F1000W filter in the MIDIS/JADES Hubble Ultra Deep Field

### v1

@@ Abstract
We report the discovery of Cerberus, an extremely red object detected with the MIRI Deep Imaging Survey (MIDIS) observations in the F1000W filter of the Hubble Ultra Deep Field. The object is detected at , with mag, and it is extremely faint in both the NIRCam data gathered by the JWST Advanced Deep Extragalactic Survey, JADES, with 30.5 mag upper limits in individual bands, as well as in the MIDIS F560W ultra deep data (29 mag, ). Analyzing the spectral energy distribution built with individual (low ) optical-to-mid-infrared filters and () stacks, we discuss the possible nature of this red NIRCam-dark source using a battery of codes. We discard the possibility of Cerberus being a Solar System body based on the proper motion in the 1-year apart JADES and MIDIS observations. A sub-stellar Galactic nature is deemed unlikely, given that the Cerberus’ relatively flat NIRCam-to-NIRCam and very red NIRCam-to-MIRI flux ratios are not consistent with any brown dwarf model. The extragalactic nature of Cerberus offers 3 possibilities: (1) A galaxy with strong emission from polycyclic aromatic hydrocarbons; the very low inferred stellar mass, M, makes this possibility highly improbable. (2) A dusty galaxy at with an inferred stellar mass M. (3) A galaxy with observational properties similar to those of the reddest little red dots discovered around , but Cerberus lying at , presenting a spectral energy distribution in the rest-frame optical dominated by emission from a dusty torus or a dusty starburst.
@@ 1 Introduction
With the development of very sensitive instruments in space telescopes such as Spitzer, the search was moved to longer wavelengths in the mid-infrared, at first compared with optical datasets (Wilson et al., 2004; Yan et al., 2004).
Remarkably, when the Hubble Space Telescope gathered extremely deep data taken in the near-infrared up to 1.6 m, with surveys such as CANDELS or the WFC3 Hubble Ultra Deep Field (Koekemoer et al., 2011; Grogin et al., 2011; Koekemoer et al., 2013), Spitzer/IRAC data were still shown to detect unique distinctively red objects which were only visible at wavelengths longer than 3 m, or at least were extremely faint in the near-infrared and optical spectral ranges. These so-called HST-dark galaxies were studied for around a decade, with significant uncertainties about their nature due to the strong sensitivity limitations of our multi-wavelength telescopes, even those operating in very different observational windows such as the (sub-)millimiter range (see, e.g., Franco et al., 2018; Alcalde Pampliega et al., 2019; Wang et al., 2019). JWST has started, from the earliest phases of the mission, to shed light on those HST-dark galaxies, i.e., mid-infrared bright near-infrared/optically faint sources which now constitute the bulk of the JWST galaxy exploration (Pérez-González et al., 2023a; Barrufet et al., 2023; Williams et al., 2023a; Gillman et al., 2023).
@@
Remarkably, with the advent of JWST, both industries, searching for red objects and for those with higher and higher redshifts, have been merged at a profound level, with the possibility of finding double-break (Lyman and Balmer) red galaxies, now known as little red dots (Labbé et al., 2023a; Matthee et al., 2023; Barro et al., 2023; Pérez-González et al., 2024), or high redshift galaxies with very high equivalent width emission lines, first discovered with Spitzer (Smit et al., 2014), now appearing in many JWST works (Pérez-González et al., 2023a; Rinaldi et al., 2023). Interestingly, some of these sources have been confirmed to present broad emission line components possibly linked to an active galactic nucleus (AGN; Kocevski et al. 2023; Greene et al. 2023).
@@ 2.1 MIRI data
MIRI data in the F1000W filter were taken as part of the MIRI Deep Imaging Survey (MIDIS, Östlin et al. 2024, in prep.) in December 2023. The observation consisted of 11 exposures, each with 100 groups, FASTR1 readout, and 10 integrations, for a total on source exposure time of 30800 s, centered on the Hubble Ultra Deep Field (HUDF). The dithering pattern was set to large-size cycling, with the 11 exposures taken in different positions on the sky separated by up to 10″. We remark that the dithering pattern, with no repeated positions, is very important for the detection of faint sources, such as the one presented here, and their distinction from detector artifacts.
The reduction of the MIRI data followed the methodology described in Pérez-González et al. (2024). Briefly, a super-background strategy is used to build background maps for each single image using all the other exposures (since they were taken during the same campaign), which results in a very homogeneous background (in terms of level and noise). Known sources are masked to avoid biasing the determination of the very local background in the super-background frame. Our F1000W image, reduced with a 60 milliarcsec pixels, presents an average 5 depth of 26.8 mag, measured in an r=04 circular aperture, and taking into account a 30% aperture correction as well as noise correlation.
@@ 2.2.1 Other MIRI data
For this letter, we combined the recently acquired MIDIS F1000W data with the F560W ultra-deep observations also carried out by MIDIS in December 2022. These data are described in Östlin et al. (2024, in prep.) and consist of hours on source taken in the HUDF, which allowed us to reach 28.7 mag 5 for point-like sources measured in an r=023 circular aperture.
@@ 2.2.2 NIRCam data
The MIRI/F1000W data were also complemented with NIRCam imaging taken by the JWST Advanced Deep Extragalactic Survey, JADES (Eisenstein et al., 2023a), Data Release 2 (Eisenstein et al. 2023b), which includes also observations from the JWST Extragalactic Medium-band Survey (JEMS, Williams et al. 2023b) and the First Reionization Epoch Spectroscopically Complete Observations (FRESCO, Oesch et al. 2023). This dataset provides a total of 14 bands from 0.9 to 4.8 m (6 at short-wavelength, SW, and 8 at long-wavelength, LW), with 5 depths ranging from 30.5 to 30.9 mag (measured in a 02 radius circular apertures). We remark that JADES is one of the deepest NIRCam surveys on the sky, only matched in depth (in some bands) by the MIDIS NIRCam-parallel project (Pérez-González et al., 2023b) and The Next Generation Deep Extragalactic Exploratory Public Near-Infrared Slitless Survey, NGDEEP (Bagley et al., 2023), which means that our search for NIRCam-dark galaxies is really probing the current depth limitations of JWST extragalactic surveys.
@@ 2.2.3 ALMA data
The Atacama Large Millimeter/submillimeter Array (ALMA) Spectroscopic Survey in the Hubble Ultra Deep Field (ASPECS) is a Cycle 4 Large Program over a 4.6 arcmin scan at 1.2 mm (Decarli et al., 2020; González-López et al., 2020) and 3.0 mm (Decarli et al., 2019; González-López et al., 2019). By covering a broader area and offering complementary data on spectral lines and redshifts, these surveys have progressively enhanced our understanding of the cosmic molecular gas density up to a redshift of approximately . The ultra-deep 1.2 mm data reaches an rms sensitivity of 9.3 Jy/beam, with beam dimensions of . The 3.0 mm data reaches 1.4 Jy/beam, with a beam.
@@ 2.2.4 MUSE data
The HUDF has been extensively studied in the last decade with the Multi Unit Spectroscopic Explorer (MUSE, Bacon et al., 2010) mounted on the Very Large Telescope (VLT) as part of the MOSAIC and UDF-10 fields (GTO programs 094.A- 0289(B), 095.A-0010(A), 096.A-0045(A) and 096.A- 0045(B), PI: R. Bacon) and the most recent MXDF observations (GTO Large Program 1101.A-0127, PI: R. Bacon). More details about these programs can be found in Bacon et al. (2017, 2023).
Briefly, MOSAIC, UDF-10, and MXDF covered the Hubble eXtreme Deep Field (XDF) within HUDF with MUSE observations in Wide-Field Mode (WFM) with each single pointing covering an area of about 1 arcmin, with a spectral wavelength range between 4700 - 9300 Å and a spectral resolving power (R) that varies from 1770 (4800 Å) to 3590 (9300 Å). The UDF-10 and MOSAIC programs were carried out without employing the ground-layer adaptive optics mode provided by the VLT Adaptive Optics Facility (AOF) and its GALACSI adaptive optics module, as detailed in Kolb et al. (2016) and Madec et al. (2018). In contrast, the MXDF program leveraged the capabilities of VLT’s AOF along with the GALACSI module to enhance its observations.
By using the WFM mode, MUSE achieves a spatial sampling of 02 02. However, the spatial resolution experienced notable variation across the different programs, with a median value of approximately 08 for the MOSAIC and UDF-10 programs, improving to 04 for the MXDF observations.
In the XDF region, which is covered by the MIDIS field, observations have achieved depths exceeding 140 hours, although the distribution of the depth in flux is not uniform across the area.
For this letter, we used the fully reduced MUSE datacubes for the MOSAIC + UDF-10 and MXDF programs, as well as the catalog of detected sources and corresponding redshifts that can be found in Bacon et al. (2023).
@@ 3 Selection and validation of NIRCam-dark sources
This letter presents a source that was discovered on the MIDIS F1000W image and found to be missed by published catalogs (Rieke et al., 2023; Eisenstein et al., 2023b). In this section, we describe the selection procedure for such types of NIRCam-dark sources.
@@ 3.1 Selection of NIRCam-dark sources
We meticulously examined each of the potential NIRCam-dark sources in the final mosaic to rule out the presence of an unflagged cosmic ray or any kind of hot and/or bad pixels that could have caused a spurious detection. We remark here that we have utilized 11 different exposures in the MIDIS F1000W survey, so the final mosaic should be free of most of these artifacts if the outlier rejection of the jwst pipeline works properly.
@@
Visual inspection of the individual exposures revealed that for 16 of the 17 candidates, there was indeed a hot pixel in one of the 11 exposures which had survived the stacking process performed to produce the final mosaic (i.e., stage 3 of the official pipeline). We automated and generalized the source validation methodology by producing stacked mosaics with a limited number of exposures, excluding 1, 2, or 3 random frames. This procedure, obviously, affects the depth. Explaining the latter case in more detail as an example, we started from our collection of 11 *_cal.fits files, and we created the 165 unique mosaics which can be obtained by excluding sets of 3 *_cal.fits files for each partial mosaic. After building the 165 realizations, we inspected the 17 potential NIRCam-dark sources, concluding that only one source consistently appeared in all 165 different partial mosaics. The remaining candidates were found to vanish in some of the realizations, in fact even in the mosaics built excluding one single exposure, indicating their most-likely spurious nature. To further validate these findings, we made an in-depth analysis of each *_cal.fits file, confirming the presence of unflagged hot and/or bad pixels in some of them that indeed led to false detections in a number of the realizations.
@@
One source of the 17 pre-selected candidates did not vanish in any of the partial mosaics, even when removing 3 random images. For this source, we measured photometry in a small r=025 aperture (which maximized on the partial mosaics), as well as the sky noise in a square 3″3″ region around, on all the mosaics produced by excluding 1, 2, and 3 exposures. We found consistent photometry and a degradation of the consistent with the differences in exposure time. For example, for the skip-3 mosaics, we found an average magnitude 27.050.19 mag, consistent with our measurement in the full mosaic (see Section 4). However, we found 2 contiguous exposures, numbers 7 and 8, which, if removed, produced a dimming of the source to 27.4 mag, still within 2 of the measurement using the full dataset. We examined these exposures in detail and found no signs of a cosmic ray hit, but a slightly enhanced signal in an extended region which could be due to the general properties of the detector (which consistently showed a dimmer region in the super-background frames for all images) or a cosmic ray shower. Given that the two relatively (and slightly) brighter images were contiguous, we do not rule out transient behavior of the source or of some detector effect. Based on the persistence of the signal from the source in all partial stacks of the MIRI F1000W dataset, we concluded that it was not spurious.
@@
The NIRCam-dark source we identified with this methodology was dubbed Cerberus, to acknowledge that it is a weird beast lurking in the darkest corners of the JWST surveys and with different interpretations/heads for its nature, as will be shown in Section 4. Its coordinates (J2000) are 03:32:38.09 (hours), 27:47:11.81 (degrees). An RGB image of the source built with NIRCam and MIRI data is presented in Figure 1, jointly with postage stamps (2″ 2″) in the NIRCam/F444W, MIRI/F560W, and MIRI/F1000W filters.
@@ 3.2 Multi-wavelength validation and observational properties of the NIRCam-dark object Cerberus
In order to further validate and characterize Cerberus, we searched for detections in the datasets described in Section 2, more specifically in the ultra-deep observations (among or even the deepest on the sky) provided by JWST/NIRCam, ALMA, and VLT/MUSE.
@@
Directly linked to our selection criteria, the NIRCam-dark Cerberus object was not present in any of the catalogs of the JADES NIRCam survey (Rieke et al., 2023; Eisenstein et al., 2023b). We then proceeded to force photometric measurements in all bands (described in Section 4) and stacked several sets of images to probe any possible low-level signal present in the images. Figure 2 provides postage stamps of the source in all JADES NIRCam bands as well as in the stacks.
The 2 rows of greyscale stamps (one presenting direct images, another one showing data convolved with a 2-pixel radius tophat filter to enhance faint signal) show no clear detection of Cerberus in any band except F460M. The brightest knot observed within a 02 radius circular aperture in that band is remarkably compact, so we decided to investigate its possible spurious nature following the same procedure as for the F1000W data, i.e., producing partial-dataset stacks. We reduced the JEMS data ourselves and found that one of the 12 images presented a single hot pixel that survived the stacking process. Indeed, in the 220 different realizations of the NIRCam/F460M dataset we built by excluding 3 random *_cal.fits files, the signal disappeared in all combinations excluding this frame. We decided to exclude this band from further analysis of Cerberus, since the central signal was also found to present an unusual shape. For the rest of the bands, we ran SExtractor using a more aggressive approach than what was used for the MIRI/F1000W data. Notwithstanding this, we were not able to detect anything at that position, except in the F460M band.
The 2 rows of postage stamps in Figure 2 shown with a red-palette and with our source marked with a blue circle depict incremental stacks constructed by averaging NIRCam bands starting from the reddest (convolved with a tophat kernel). The image on the left-end of the first row of these stamps combines F480M and F460M, the second adds F444W to the previous, the next one adds F430M, etc., and so forth. The final image adds all NIRCam bands from F182M to F444W. The second row of red stamps excludes F460M from the stacks, given the confirmed spurious nature of part of the signal in that filter. When we ran SExtractor, adopting the same aggressive method as for the individual NIRCam bands, we actually detected the source precisely at the location of our initial MIRI-detected NIRCam-dark candidate selected at 10 m. The detection presents and it is compatible with a point-like source (but see discussion about morphology below). This test definitively demonstrated the existence of our MIRI-selected NIRCam-dark source, showing that this elusive source has a counterpart observable in the NIRCam bands exclusively by stacking very deep data, i.e., it lies near the detection limit of the JADES dataset (which is among or possibly even the deepest on the sky, jointly with MIDIS and NGDEEP, as we previously remarked).
The row presenting blue-palette stamps in Figure 2 shows stacks of NIRCam images starting from the bluest and adding one filter at a time. The image on the left-end of the row corresponds to F090WF115W, then F150W is added, etc, and the last image includes all SW filters up to F210M. The source is much fainter in these blue stacks, some signal at the level is tentatively detected in the last stack (see Section 4).
Apart from the incremental NIRCam stacks, in the following section we will use four sets of disjoint stacks, combining 2, 3, 4 and 5 bands, always starting from the reddest and not repeating them in different stacks of the same set. These stacks provided detections in the 2-band set, specifically in the F480MF444W (30.0 mag) and F356WF335M (30.6 mag) combinations, and in one 3-band stack, the one combining F410MF356WF335M (30.5 mag).
@@
Figure 3 shows a stacked NIRCam image (including all filters at m) convolved to the resolution of the F1000W data (FWHM=0328). Each calibrated individual NIRCam filter was rebinned to a 006 pixel scale and PSF-matched to F1000W using WebbPSF (v. 1.2.1 Perrin et al., 2014) models and the method described in Melinder et al. (2023), and then median-stacked. We measured the centroid of the source for both JWST instruments (using photutils with a circular footprint of 3 pixels radius), finding an agreement within 0.26 pixels, i.e., 0016. Remarkably, the source is slightly extended to the NW in both images. The morphology of the source is similar for the NIRCam stack and the MIRI F1000W image, with a tadpole-like shape extending to the upper right, further strengthening the positive detection of the source with NIRCam. We also notice a qualitative resemblance between the structure of Cerberus and GN-z11 (Oesch et al., 2016; Tacchella et al., 2023), with a concentrated flux distribution identified with the core of the galaxy and an extended haze component.

### v2

@@ Abstract
We report the discovery of Cerberus, an extremely red object detected with the MIRI Deep Imaging Survey (MIDIS) observations in the F1000W filter of the Hubble Ultra Deep Field. The object is detected at , with mag, and undetected in the NIRCam data gathered by the JWST Advanced Deep Extragalactic Survey, JADES, fainter than the 30.0-30.5 mag detection limits in individual bands, as well as in the MIDIS F560W ultra-deep data (29 mag, ). Analyzing the spectral energy distribution built with low- () measurements in individual optical-to-mid-infrared filters and higher () in stacked NIRCam data, we discuss the possible nature of this red NIRCam-dark source using a battery of codes. We discard the possibility of Cerberus being a Solar System body based on the proper motion in the 1-year apart JADES and MIDIS observations. A sub-stellar Galactic nature is deemed unlikely, given that the Cerberus’ relatively flat NIRCam-to-NIRCam and very red NIRCam-to-MIRI flux ratios are not consistent with any brown dwarf model. The extragalactic nature of Cerberus offers 3 possibilities: (1) A galaxy with strong emission from polycyclic aromatic hydrocarbons; the very low inferred stellar mass, M⊙, makes this possibility highly improbable. (2) A dusty galaxy at with an inferred stellar mass M⊙. (3) A galaxy with observational properties similar to those of the reddest little red dots discovered around , but Cerberus lying at , with the rest-frame optical dominated by emission from a dusty torus or a dusty starburst.
@@ 1 Introduction
With the development of very sensitive instruments onboard space telescopes such as Spitzer, the search was moved to longer wavelengths in the mid-infrared, at first compared with optical datasets (Wilson et al., 2004; Yan et al., 2004).
Remarkably, when the Hubble Space Telescope gathered extremely deep data taken in the near-infrared up to 1.6 m, with surveys such as CANDELS or the WFC3 Hubble Ultra Deep Field (Koekemoer et al., 2011; Grogin et al., 2011; Koekemoer et al., 2013), Spitzer/IRAC data were still shown to detect unique distinctively red objects which were only visible at wavelengths longer than 3 m, or at least were extremely faint in the near-infrared and optical spectral ranges. These so-called HST-dark galaxies were studied for around a decade, with significant uncertainties about their nature due to the strong sensitivity limitations of our multi-wavelength telescopes, even those operating in very different observational windows such as the (sub-)millimeter range (see, e.g., Simpson et al., 2014; Franco et al., 2018; Alcalde Pampliega et al., 2019; Wang et al., 2019). JWST has started, from the earliest phases of the mission, to shed light on those HST-dark galaxies, i.e., mid-infrared bright near-infrared/optically faint sources which now constitute the bulk of the JWST galaxy exploration (Pérez-González et al., 2023a; Nelson et al., 2023; Barrufet et al., 2023; Williams et al., 2023a; Gillman et al., 2023).
@@
Remarkably, with the advent of JWST, both industries, searching for red objects and for those with higher and higher redshifts, have been merged at a profound level, with the possibility of finding double-break (Lyman and Balmer) red galaxies (Labbé et al., 2023a), now known as little red dots (Furtak et al., 2023; Übler et al., 2023; Matthee et al., 2023; Barro et al., 2023; Greene et al., 2023; Williams et al., 2023a; Killi et al., 2023; Pérez-González et al., 2024; Kokorev et al., 2024), or high redshift galaxies with very high equivalent width emission lines, first discovered with Spitzer (Smit et al., 2014), now appearing in many JWST works (Pérez-González et al., 2023a; Rinaldi et al., 2023). Interestingly, some of these sources have been confirmed to present broad emission line components possibly linked to an active galactic nucleus (AGN; Kocevski et al. 2023; Greene et al. 2023).
@@ 2.1 MIRI data
MIRI data in the F1000W filter were taken as part of the MIRI Deep Imaging Survey (MIDIS, Östlin et al. 2024, in prep.; PID 1283) in December 2023. The observation consisted of 11 exposures, each with 100 groups, FASTR1 readout, and 10 integrations, for a total on source exposure time of 30800 s, centered on the Hubble Ultra Deep Field (HUDF). The dithering pattern was set to large-size cycling, with the 11 exposures taken in different positions on the sky separated by up to 10″.
The reduction of the MIRI data followed the methodology described in Pérez-González et al. (2024), based on the official jwst pipeline (v1.13.4, jwst_1202.pmap) and offline bespoke procedures. Briefly, a super-background strategy is used to build background maps for each single image using all the other exposures (since they were taken during the same campaign), which results in a very homogeneous background in terms of level and noise. Quantitatively for the data used in this letter, the standard deviation of the background pixels for each cal.fits file obtained with the standard official pipeline is reduced by a factor of by using our super-background method, and the depth in the final mosaic improves by 0.8 mag. We refer the reader to Appendix A in Pérez-González et al. (2024) for more details. Known sources are masked to avoid biasing the determination of the very local background in the super-background image. Our F1000W final mosaic, reduced with a 60 milliarcsec pixels, presents an average 5 depth of 26.8 mag, measured in a 04 radius circular aperture, and taking into account noise correlation (linked to the drizzling method used in the mosaicking) and after applying the aperture correction (a 1.3 factor).
@@ 2.2.1 Other MIRI data
For this letter, we combined the recently acquired MIDIS F1000W data with the F560W ultra-deep observations also carried out by MIDIS in December 2022. These data will be described in Östlin et al. (2024, in prep.) and consist of hours on source taken in the HUDF, which allowed us to reach 28.7 mag 5 for point-like sources measured in an r=023 circular aperture (the radius being chosen to ensure a % encircled energy).
@@ 2.2.2 NIRCam data
The MIRI/F1000W data were also complemented with NIRCam imaging taken by the JWST Advanced Deep Extragalactic Survey, JADES (Eisenstein et al., 2023a), Data Release 2 (Eisenstein et al. 2023b), which includes observations from the JWST Extragalactic Medium-band Survey (JEMS, Williams et al. 2023b) and the First Reionization Epoch Spectroscopically Complete Observations (FRESCO, Oesch et al. 2023). This dataset provides a total of 14 bands from 0.9 to 4.8 m (6 at short-wavelength, SW, and 8 at long-wavelength, LW), with 5 depths ranging from 30.5 to 30.9 mag (measured in a 02 radius circular apertures). We remark that JADES is one of the deepest NIRCam surveys on the sky, only matched in depth (in some bands) by the MIDIS NIRCam-parallel project (Pérez-González et al., 2023b) and The Next Generation Deep Extragalactic Exploratory Public Near-Infrared Slitless Survey, NGDEEP (Bagley et al., 2023), which means that our search for NIRCam-dark galaxies is really probing the current depth limitations of JWST extragalactic surveys.
@@ 2.2.3 ALMA data
The Atacama Large Millimeter/submillimeter Array (ALMA) Spectroscopic Survey in the Hubble Ultra Deep Field (ASPECS) is a Cycle 4 Large Program over a 4.6 arcmin2 scan at 1.2 mm (Decarli et al., 2020; González-López et al., 2020) and 3.0 mm (Decarli et al., 2019; González-López et al., 2019). The ultra-deep 1.2 mm data reaches an rms sensitivity of 9.3 Jy/beam, with beam dimensions of . The 3.0 mm data reaches 1.4 Jy/beam, with a beam.
@@ 2.2.4 MUSE data
The HUDF has been extensively studied in the last decade with the Multi Unit Spectroscopic Explorer (MUSE, Bacon et al., 2010) mounted on the Very Large Telescope (VLT) as part of the MOSAIC and UDF-10 fields (GTO programs 094.A-0289(B), 095.A-0010(A), 096.A-0045(A) and 096.A-0045(B), PI: R. Bacon) and the most recent MXDF observations (GTO Large Program 1101.A-0127, PI: R. Bacon). The MUSE data has a spectral wavelength range between 4700 – 9300 Å and a spectral resolving power (R) that ranges from 1770 (4800 Å) to 3590 (9300 Å). The HUDF dataset was taken during 140 hours on-source with the VLT Adaptive Optics Facility (AOF) and its GALACSI adaptive optics module, as detailed in Kolb et al. (2016) and Madec et al. (2018), achieving a spatial sampling of 02 02, with strong variations of up to a factor of . More details about these programs can be found in Bacon et al. (2017, 2023).
@@ 3 Selection and validation of NIRCam-dark sources
This letter presents a source that was discovered on the MIDIS F1000W image and found not to be included in existing NIRCam published catalogs (Rieke et al., 2023; Eisenstein et al., 2023b). In this section, we describe the selection procedure for such types of NIRCam-dark sources.
@@ 3.1 Selection of NIRCam-dark sources
We meticulously examined each of the potential NIRCam-dark sources in the final mosaic to rule out the presence of an unflagged cosmic ray or any kind of hot and/or bad pixels that could have caused a spurious detection.
@@
Visual inspection of the individual exposures revealed that for 16 of the 17 candidates, there was indeed a hot pixel in one of the 11 exposures which had survived the flagging and stacking process performed to produce the final mosaic. We automated and generalized the source validation methodology by producing stacked mosaics with a limited number of exposures, excluding 1, 2, or 3 of the full set randomly. This procedure, obviously, affects the depth. Explaining the latter case in more detail as an example, we started from our collection of 11 *_cal.fits files, and we created the 165 unique mosaics which can be obtained by excluding sets of 3 *_cal.fits files for each partial mosaic. After building the 165 realizations, we inspected the 17 potential NIRCam-dark sources, concluding that only one source consistently appeared in all 165 different partial mosaics. The remaining candidates were found to vanish in some of the realizations, in fact even in the mosaics built excluding one single exposure, indicating their most-likely spurious nature. To further validate these findings, we made an in-depth analysis of each *_cal.fits file, confirming the presence of unflagged hot and/or bad pixels in some of them that indeed led to false detections in a number of the realizations.
@@
One source of the 17 pre-selected candidates did not vanish in any of the partial mosaics, even when removing 3 random images. For this source, we measured photometry in a small 025 radius aperture (which maximized on the partial mosaics), as well as the sky noise in a square 3″3″ region around, on all the mosaics produced by excluding 1, 2, and 3 exposures. We found consistent photometry and a degradation of the consistent with the differences in exposure time. For example, for the skip-3 mosaics, we found an average magnitude 27.050.19 mag, consistent with our measurement in the full mosaic (see Section 4).
However, we found 2 contiguous exposures, numbers 7 and 8, which, if removed, produced a dimming of the source to 27.4 mag, still within 2 of the measurement using the full dataset. Each one of these exposures has an integration time of 2800 s, with the start of the first exposure and the end of the second separated by 5700 s. We examined these exposures in detail and found no signs of a cosmic ray hit, but a slightly enhanced signal in an extended region. This could be due to the general properties of the detector (which consistently showed a dimmer region in the super-background frames for all images). It could also be due to a cosmic ray shower, which are known to affect MIRI imaging observations (Dicken et al., 2024), with an incidence of several per exposure. Cosmic ray showers are extended persistence artifacts from cosmic ray hits and typically affect the detector for several minutes (Wright et al., 2023). However, a high flux cosmic ray hit could cause a shower lasting long enough to enhance the signal seen in these exposures. We also note that the enhancement in signal is stronger in exposure 7 compared to 8, which is also consistent with persistence (see Álvarez-Márquez et al., 2023). Given that the two relatively (and slightly) brighter images were contiguous, we do not rule out a transient behavior of the source. Based on the perseverance of the signal from the source in all partial stacks of the MIRI F1000W dataset, we concluded that it was not spurious.
@@
The NIRCam-dark source we identified with this methodology was dubbed Cerberus, to acknowledge that it is a weird beast lurking in the darkest corners of the JWST surveys and with different interpretations/heads for its nature, as will be shown in Section 4. Its coordinates (J2000) are 03:32:38.09 (hours), 27:47:11.81 (degrees). NIRCam and MIRI images of Cerberus are presented in Figure 1.
@@ 3.2 Multi-wavelength validation and observational properties of the NIRCam-dark object Cerberus
In order to further validate and characterize Cerberus, we searched for detections by forcing photometric measurements in the datasets described in Section 2, more specifically in the ultra-deep observations (among or even the deepest on the sky) provided by JWST/NIRCam, ALMA, and VLT/MUSE.
3.2.1 JWST/NIRCam measurements
@@
Directly linked to our selection criteria, the NIRCam-dark Cerberus object was not present in any of the catalogs of the JADES NIRCam survey (Rieke et al., 2023; Eisenstein et al., 2023b). We then proceeded to force photometric measurements in all bands (described in Section 4). Forced photometric measurements in all NIRCam individual bands do not provide any measurement signal-to-noise ratio .
We then stacked NIRCam images to search for faint emission. We constructed two different sets of stacks. One set consisted in cumulative stacks constructed by averaging NIRCam bands starting from the reddest and from the bluest. Another set consisted of disjoint stacks (to avoid noise spikes in one band affecting the measurements in cumulative stacks), combining 2, 3, 4 and 5 bands, always starting from the reddest and not repeating them in different stacks of the same set.
@@
Figure 2 shows a stacked NIRCam image (including all filters at m) convolved to the resolution of the F1000W data (FWHM=0328). Each calibrated individual NIRCam filter was rebinned to a 006 pixel scale and PSF-matched to F1000W using WebbPSF (v. 1.2.1 Perrin et al., 2014) models and the method described in Melinder et al. (2023), and then median-stacked. We measured the centroid of the source for both JWST instruments (using photutils with a circular footprint of 3 pixels radius), finding an agreement within 0.26 pixels, i.e., 0016. Remarkably, the source is slightly extended to the NW in both images. The morphology of the source is similar for the NIRCam stack and the MIRI F1000W image, with a tadpole-like shape extending to the upper right, confirming a detection of the source with NIRCam stacked data (see next section). We also notice a qualitative resemblance between the structure of Cerberus and the galaxy GN-z11 (Oesch et al., 2016; Tacchella et al., 2023), with a concentrated flux distribution identified with the core of the galaxy and an extended haze component.
3.2.2 MUSE measurements